# Recursive regularised lattice Boltzmann method for magnetohydrodynamics

### v1

@@ I Introduction
An alternative stabilisation strategy is provided by regularised lattice Boltzmann methods, in which the non-equilibrium distribution is reconstructed from a truncated Hermite expansion malaspinas2015increasing. More recently, recursive regularisation (RR) techniques have been proposed to systematically reconstruct higher-order non-equilibrium moments from lower-order information, leading to improved isotropy and robustness in hydrodynamic flows coreixas2017recursive; coreixas2020highorder. Despite these advances, including extensions to multiphase flows LU2026105500, recursive regularisation has not yet been systematically explored in the context of magnetohydrodynamic lattice Boltzmann models. Existing work on its application to MHD remains limited in scope and does not provide a comprehensive assessment 10.1007/978-981-95-3401-2_35. The present study fills this gap by offering the first systematic investigation of the method for MHD flows, rigorously assessing accuracy, stability, and scalability over a wide range of regimes. Indeed, we aim to assess whether the stabilisation and isotropy benefits of recursive regularisation observed in hydrodynamic flows carry over to magnetohydrodynamic systems.
@@ II.2 Recursive regularised lattice Boltzmann method
$a_{1,\alpha\beta}^{(2)}$ $=-\nu\rho c_{s}^{2}\left[\partial_{\alpha}u_{\beta}+\partial_{\beta}u_{\alpha}-(\partial_{\gamma}u_{\gamma})\delta_{\alpha\beta}\right],$
@@ III.1 Low-Reynolds-number flow
The sequence shown in Fig. 1 illustrates the temporal evolution of the magnetic field obtained by our present RR procedure. Initially, the field retains the smooth, large-scale structure imposed by the initial condition, with well-organised alternating regions of magnetic intensity. As non-linear coupling intensifies, magnetic field lines undergo progressive stretching and folding, leading to the formation of thin filaments and elongated gradients. These structures correspond to the emergence of current sheets aligned with evolving shear layers, signalling the onset of magnetic reconnection. At later times, the magnetic topology becomes increasingly intermittent and complex, accompanied by a breakdown of the initial symmetries. This evolution directly underlies the growth of $j_{\max}$ and $\omega_{\max}$ observed in the tables and reflects the transition toward fully non-linear MHD dynamics. Moreover, it corroborates the observations previously obtained by other collision operators dellar2002lbm_mhd; DeRosis03062018; DeRosis2017PRE; 10.1063/5.0058884.
@@
Figure 1: Low-Reynolds-number flow: evolution of the magnetic field at selected time instants obtained by the RR LBM. The sequence illustrates the progressive distortion of the initial large-scale structures, the formation of thin current sheets, and the emergence of a highly intermittent magnetic field as the flow transitions toward fully non-linear MHD dynamics.
@@ III.2 Turbulent regimes
We next investigate strongly non-linear regimes corresponding to $\mathrm{Re}=2500$ and $5000$. Tables 47 report the evolution of the peak current density and vorticity across progressively refined grids. These quantities directly probe the formation of thin current sheets and intense shear layers associated with magnetic reconnection and the ensuing turbulent cascade.
@@
Figure 3: Temporal evolution of the peak electric current density $j_{\max}=\max_{\Omega}|j_{z}|$ for the Orszag–Tang vortex at $\mathrm{Re}=200\pi$ (magenta solid line), $\mathrm{Re}=2500$ (brown dashed line), and $\mathrm{Re}=5000$ (olive line with smaller dashes). The initial growth reflects large-scale magnetic field-line stretching, followed by a rapid intensification associated with current-sheet thinning and the onset of magnetic reconnection. Increasing Reynolds number leads to stronger current concentration and sustained intermittency, characteristic of turbulent MHD dynamics. Findings are obtained by the present RR scheme.
@@
Figure 4: Time evolution of kinetic (solid lines) and magnetic (dashed lines) energies for the Orszag–Tang vortex at $\mathrm{Re}=200\pi$ (orange), $\mathrm{Re}=2500$ (navy), and $\mathrm{Re}=5000$ (dark green). At early times, magnetic energy dominates as the initial field is stretched by the flow. As non-linear interactions intensify, magnetic energy is rapidly converted into kinetic energy through current-sheet formation and magnetic reconnection. Increasing Reynolds number leads to stronger energy transfer, delayed saturation, and enhanced turbulent mixing, with higher-Re cases exhibiting sustained kinetic activity and slower magnetic dissipation characteristic of fully developed MHD turbulence. Findings are obtained by the present RR scheme.
@@ III.3 Computational cost
Across the entire range of resolutions considered, the RR model displays the highest runtime. This behaviour is a direct consequence of the more elaborate collision step, which involves the projection of non-equilibrium populations onto Hermite polynomials followed by recursive reconstruction procedures. While these operations improve numerical stability, they introduce a measurable computational overhead that persists as the problem size increases.
@@
We can conclude that moment-based collision operators provide a favourable balance between computational efficiency and modelling sophistication, while the RR approach trades increased runtime for enhanced numerical robustness.
@@ IV Conclusions
The Orszag–Tang vortex was used as a stringent benchmark to quantify accuracy, dissipation, and robustness against established collision models (BGK, raw-moment MRT, and central-moment formulations) under otherwise identical settings. In the well-resolved early evolution, all schemes closely match spectral reference data for peak current density and vorticity. As the flow develops sharp gradients and current sheets, the recursive regularisation introduces a modest additional dissipation on coarse grids, leading to slightly lower extrema compared with BGK. Importantly, this discrepancy diminishes under grid refinement: on the finest meshes, the RR results converge to the same level of agreement achieved by the competing collision operators, confirming consistency with the reference solution in the physically relevant regime.
We further examined the solenoidal constraint on the magnetic field. All tested lattice Boltzmann formulations display clear rapid decay of the maximum divergence error with increasing resolution, and the differences between collision models remain negligible. This indicates that the divergence behaviour is primarily controlled by the discrete magnetic-field representation (D2Q5) rather than by the regularisation of the fluid non-equilibrium moments. Qualitatively, the RR scheme reproduces the canonical Orszag–Tang dynamics, including the progressive distortion of large-scale magnetic structures, the formation of thin current sheets, and the emergence of intermittent filamentary features, in line with established MHD studies within the LBM community.
@@
From a computational standpoint, the recursive regularisation entails a moderate increase in per-time-step cost due to the additional Hermite projections and recursive reconstructions performed during the collision step. Nevertheless, the algorithm preserves the strictly local character of the lattice Boltzmann method and maintains linear scaling with system size. As a result, the added overhead remains controlled and does not compromise the favourable parallel efficiency or memory-access patterns of standard LBM formulations. In practical terms, the improved numerical robustness offered by the RR scheme can offset this additional cost by enabling stable simulations at coarser resolutions or under more demanding physical conditions.

### v2

@@ I Introduction
An alternative stabilisation strategy is provided by regularised lattice Boltzmann methods, in which the non-equilibrium distribution is reconstructed from a truncated Hermite expansion malaspinas2015increasing. More recently, recursive regularisation ($R$) techniques have been proposed to systematically reconstruct higher-order non-equilibrium moments from lower-order information, leading to improved isotropy and robustness in hydrodynamic flows coreixas2017recursive; coreixas2020highorder. Despite these advances, including extensions to multiphase flows LU2026105500, recursive regularisation has not yet been systematically explored in the context of magnetohydrodynamic lattice Boltzmann models. Existing work on its application to MHD remains limited in scope and does not provide a comprehensive assessment 10.1007/978-981-95-3401-2_35. The present study fills this gap by offering the first systematic investigation of the method for MHD flows, rigorously assessing accuracy, stability, and scalability over a wide range of regimes. Indeed, we aim to assess whether the stabilisation and isotropy benefits of recursive regularisation observed in hydrodynamic flows carry over to magnetohydrodynamic systems.
@@ II.2 Recursive regularised lattice Boltzmann method
$a_{1,\alpha\beta}^{(2)}$ $=-\nu\rho c_{s}^{2}\left[\partial_{\alpha}u_{\beta}+\partial_{\beta}u_{\alpha}-(\partial_{\gamma}u_{\gamma})\delta_{\alpha\beta}\right]-M_{\alpha\beta},$
@@ III.1 Low-Reynolds-number flow
The sequence shown in Fig. 1 illustrates the temporal evolution of the magnetic field obtained by our present $R$ procedure. Initially, the field retains the smooth, large-scale structure imposed by the initial condition, with well-organised alternating regions of magnetic intensity. As non-linear coupling intensifies, magnetic field lines undergo progressive stretching and folding, leading to the formation of thin filaments and elongated gradients. These structures correspond to the emergence of current sheets aligned with evolving shear layers, signalling the onset of magnetic reconnection. At later times, the magnetic topology becomes increasingly intermittent and complex, accompanied by a breakdown of the initial symmetries. This evolution directly underlies the growth of $j_{\max}$ and $\omega_{\max}$ observed in the tables and reflects the transition toward fully non-linear MHD dynamics. Moreover, it corroborates the observations previously obtained by other collision operators dellar2002lbm_mhd; DeRosis03062018; DeRosis2017PRE; 10.1063/5.0058884.
@@
Figure 1: Low-Reynolds-number flow: evolution of the magnetic field at selected time instants obtained by the $R$ LBM. The sequence illustrates the progressive distortion of the initial large-scale structures, the formation of thin current sheets, and the emergence of a highly intermittent magnetic field as the flow transitions toward fully non-linear MHD dynamics.
@@ III.2 Turbulent regimes
We next investigate strongly non-linear regimes corresponding to $\mathrm{Re}=2500$ and 50005000. Tables 47 report the evolution of the peak current density and vorticity across progressively refined grids. These quantities directly probe the formation of thin current sheets and intense shear layers associated with magnetic reconnection and the ensuing turbulent cascade.
@@
Figure 3: Temporal evolution of the peak electric current density $j_{\max}=\max_{\Omega}|j_{z}|$ for the Orszag–Tang vortex at $\mathrm{Re}=200\pi$ (magenta solid line), $\mathrm{Re}=2500$ (brown dashed line), and $\mathrm{Re}=5000$ (olive line with smaller dashes). The initial growth reflects large-scale magnetic field-line stretching, followed by a rapid intensification associated with current-sheet thinning and the onset of magnetic reconnection. Increasing Reynolds number leads to stronger current concentration and sustained intermittency, characteristic of turbulent MHD dynamics. Findings are obtained by the present $R$ scheme.
@@
Figure 4: Time evolution of kinetic (solid lines) and magnetic (dashed lines) energies for the Orszag–Tang vortex at $\mathrm{Re}=200\pi$ (orange), $\mathrm{Re}=2500$ (navy), and $\mathrm{Re}=5000$ (dark green). At early times, magnetic energy dominates as the initial field is stretched by the flow. As non-linear interactions intensify, magnetic energy is rapidly converted into kinetic energy through current-sheet formation and magnetic reconnection. Increasing Reynolds number leads to stronger energy transfer, delayed saturation, and enhanced turbulent mixing, with higher-Re cases exhibiting sustained kinetic activity and slower magnetic dissipation characteristic of fully developed MHD turbulence. Findings are obtained by the present $R$ scheme.
@@ III.3 Computational cost
Across the entire range of resolutions considered, the $R$ model displays the highest runtime. This behaviour is a direct consequence of the more elaborate collision step, which involves the projection of non-equilibrium populations onto Hermite polynomials followed by recursive reconstruction procedures. While these operations improve numerical stability, they introduce a measurable computational overhead that persists as the problem size increases.
@@
We can conclude that moment-based collision operators provide a favourable balance between computational efficiency and modelling sophistication, while the $R$ approach trades increased runtime for enhanced numerical robustness.
@@ IV Conclusions
The Orszag–Tang vortex was used as a stringent benchmark to quantify accuracy, dissipation, and robustness against established collision models (BGK, raw-moment MRT, and central-moment formulations) under otherwise identical settings. In the well-resolved early evolution, all schemes closely match spectral reference data for peak current density and vorticity. As the flow develops sharp gradients and current sheets, the recursive regularisation introduces a modest additional dissipation on coarse grids, leading to slightly lower extrema compared with BGK. Importantly, this discrepancy diminishes under grid refinement: on the finest meshes, the $R$ results converge to the same level of agreement achieved by the competing collision operators, confirming consistency with the reference solution in the physically relevant regime.
We further examined the solenoidal constraint on the magnetic field. All tested lattice Boltzmann formulations display clear rapid decay of the maximum divergence error with increasing resolution, and the differences between collision models remain negligible. This indicates that the divergence behaviour is primarily controlled by the discrete magnetic-field representation (D2Q5) rather than by the regularisation of the fluid non-equilibrium moments. Qualitatively, the $R$ scheme reproduces the canonical Orszag–Tang dynamics, including the progressive distortion of large-scale magnetic structures, the formation of thin current sheets, and the emergence of intermittent filamentary features, in line with established MHD studies within the LBM community.
@@
From a computational standpoint, the recursive regularisation entails a moderate increase in per-time-step cost due to the additional Hermite projections and recursive reconstructions performed during the collision step. Nevertheless, the algorithm preserves the strictly local character of the lattice Boltzmann method and maintains linear scaling with system size. As a result, the added overhead remains controlled and does not compromise the favourable parallel efficiency or memory-access patterns of standard LBM formulations. In practical terms, the improved numerical robustness offered by the $R$ scheme can offset this additional cost by enabling stable simulations at coarser resolutions or under more demanding physical conditions.